# Segment-Aware Conditioning for Training-Free Intra-Utterance Emotion and Duration Control in Text-to-Speech

### v1

@@ Abstract
While controllable Text-to-Speech (TTS) has achieved notable progress, most existing methods remain limited to inter-utterance-level control, making fine-grained intra-utterance expression challenging due to their reliance on non-public datasets or complex multi-stage training. In this paper, we propose a training-free controllable framework for pretrained zero-shot TTS to enable intra-utterance emotion and duration expression. Specifically, we propose a segment-aware emotion conditioning strategy that combines causal masking with monotonic stream alignment filtering to isolate emotion conditioning and schedule mask transitions, enabling smooth intra-utterance emotion shifts while preserving global semantic coherence. Based on this, we further propose a segment-aware duration steering strategy to combine local duration embedding steering with global EOS logit modulation, allowing local duration adjustment while ensuring globally consistent termination. To eliminate the need for segment-level manual prompt engineering, we construct a 30,000-sample multi-emotion and duration-annotated text dataset to enable LLM-based automatic prompt construction. Extensive experiments demonstrate that our training-free method not only achieves state-of-the-art intra-utterance consistency in multi-emotion and duration control, but also maintains baseline-level speech quality of the underlying TTS model. Audio samples are available at https://aclanonymous111.github.io/TED-TTS-DemoPage/.
Segment-Aware Conditioning for Training-Free Intra-Utterance Emotion and Duration Control in Text-to-Speech
Qifan Liang1thanks: Equal contribution., Yuansen Liu [^1] footnotemark: 1, Ruixin Wei [^1] footnotemark: 1, Nan Lu1, Junchuan Zhao1, Ye Wang1thanks: Corresponding author. 1School of Computing, National University of Singapore {liangqifan,e1553665,e1538320,nlu}@u.nus.edu, {junchuan,wangye}@comp.nus.edu.sg
@@ 1 Introduction
Humans naturally regulate emotional expression and speaking pace during speech in a dynamic and flexible manner, reflecting changes in semantics, emphasis, and discourse intent. How to replicate such intra-utterance expressiveness remains a central challenge in building human-like Text-to-Speech (TTS) synthesis systems.
@@
These challenges naturally raise an important question: Is it possible to achieve stable segment-level emotion transitions and duration control without retraining the model? In this paper, as shown in Fig. 1, we revisit controllable TTS from an inference-time perspective and propose the first training-free intra-utterance emotion and duration control framework. Rather than introducing additional predictors or retraining the acoustic model, our approach focuses on restructuring how conditioning information is accessed and updated during autoregressive decoding. Specifically, for multi-emotion control, we propose a segment-aware emotion conditioning strategy that combines causal masking with monotonic stream alignment filtering algorithm, which jointly isolates segment-specific emotion conditioning and performs online text-semantic alignment to schedule mask transitions, enabling smooth intra-utterance emotion shifts while preserving global semantic coherence. To enable multi-duration control, we further propose a segment-aware duration steering strategy to incorporate local duration embedding steering with global EOS logit modulation, allowing segment-level pacing adjustment while ensuring globally consistent sequence termination. Besides, we construct a multi-emotion and duration-annotated text dataset with 30,000 samples and fine-tune Qwen3-8B to enable LLM-based automatic prompt construction, thereby eliminating the need for segment-level manual segmentation and prompt engineering. Extensive experiments demonstrate that our method achieves state-of-the-art performance in stable intra-utterance multi-emotion transitions and duration control, while preserving the strong zero-shot synthesis capability of the underlying TTS model without any additional training. Our contributions are summarized as follows:
@@
  • We propose a segment-aware duration steering strategy that achieves local segment duration control while preserving global consistent sequence termination.
@@ 3 Method
3.1 Automatic Prompt Construction
Existing intra-utterance controllable TTS systems require manual text segmentation and segment-level emotion and duration specification, which is labor-intensive in real-world scenarios. To eliminate manual prompt engineering, we fine-tune the Qwen3-8B LLM to automatically transform raw user text into structured multi-segment prompts. As a prerequisite, we construct a dedicated Multi-Emotion and Duration-annotated text dataset (MED-TTS) with 30,000 samples, which is used to supervise emotion-aware text segmentation, natural language emotion description generation, and segment-level speech duration estimation. As illustrated below, MED-TTS is synthesized using LLM through a structured pipeline consisting of generation, annotation, and verification.
@@ 3.1 Automatic Prompt Construction
Step 1: Content text generation. GPT-4o is prompted to generate emotion-rich English and Chinese texts with continuous intra-utterance emotional transitions spanning multiple phases from seven core emotions (happy, sad, angry, surprised, fearful, disgusted, and neutral), covering descriptive, dialog-style, and observational content in three text categories.
Step 2: Multi-segment prompt annotation. To enable precise segment-level control, DeepSeek-Chat is prompted to decompose each text into emotion-specific segments and assigns each segment a concise natural language emotion description and a realistic duration estimate, yielding structured emotion-duration sequences compatible with controllable TTS inputs.
Step 3: Post-processing and manual verification. Automatic checks are finally applied to filter samples with formatting errors, missing fields, or invalid segment boundaries, followed by systematic manual verification of outputs from both Step 1 and Step 2. Based on this dataset, we perform supervised fine-tuning with LoRA on the Qwen3-8B large language model, enabling automatic construction of segment-level TTS prompts without manual prompt engineering. Detailed prompting strategies, step-wise checklists, dataset statistics, and fine-tuning details are provided in the Appendix A.
@@ 3.2 Segment-Aware Emotion Conditioning
2D Causal Attention Mask. To resolve the misalignment between continuous generation and segment-level conditions, we design a 2D causal attention mask that disentangles condition visibility from semantic context. The mask preserves standard causal attention among text and semantic tokens across segment boundaries, ensuring globally coherent semantic generation while strictly restricting access to condition embeddings to be segment-local. Specifically, for any token that belongs to the $m$-th segment (either a text token in $\mathrm{X}_{m}$ or a generated semantic token that currently aligns to $\mathrm{X}_{m}$), attention is allowed to attend only to its corresponding condition embedding $\mathbf{C}_{m}$, while all other condition embeddings $\{\mathbf{C}_{j}\mid j\neq m\}$ are masked out. Meanwhile, each condition embedding $\mathbf{C}_{m}$ is prevented from attending to other condition embeddings, avoiding cross-condition information leakage. After that, as shown in the bottom of Fig. 2, emotional style is governed exclusively by the locally active condition, whereas semantic content remains globally visible through standard causal context.
@@ 5 Results and Evaluation
5.1 Comparison with Reference Models
5.2 Ablation Study
@@
Method *0.75 *0.875 *1 *1.125 *1.25
@@
Monotonic Stream Alignment Evaluation. To evaluate the effectiveness of MSA, we visualize alignment results under different settings in Fig. 4 and report the mean absolute error (MAE) of segment boundary positions. Raw attention maps exhibit diffuse and locally non-monotonic patterns, making greedy alignment highly sensitive to noise and leading to unstable trajectories and frequent segment switching failures. Introducing the monotonic stream constraint alleviates this issue and reduces MAE to 0.216, but residual attention uncertainty still causes instability. By further incorporating the observation component, MSA effectively suppresses alignment uncertainty, enforces smooth monotonic trajectories, and reduces MAE to 0.157, yielding precise emotion transitions closely aligned with the ground-truth boundaries.
Duration-Specified Evaluation. We evaluate duration-specified speech synthesis under five segment-level scaling factors (*0.75, *0.875, *1.0, *1.125, and *1.25), comparing our full system with ablated variants and an IndexTTS2 baseline. As shown in Tab. 4, our method consistently achieves the lowest semantic token number error across all settings, reducing the error by 3.53% and 2.41% on average compared to variants without local steering and global EOS control, respectively. Relative to the baseline without explicit duration control, our approach further yields a 5.07% average error reduction, demonstrating accurate and robust duration control across diverse segment-level targets.
@@ 6 Conclusion
In this paper, we propose the first training-free controllable framework to enable intra-utterance emotion and duration control in pretrained zero-shot TTS. By introducing segment-aware emotion conditioning and duration steering from an inference-time perspective, our method achieves smooth emotional transitions and temporal regulation within a single utterance. Extensive experiments demonstrate that our method not only delivers state-of-the-art intra-utterance controllability, but also preserves baseline-level speech quality of the underlying TTS model.
@@ A.1 LLM-Based Prompting Strategy
MED-TTS is synthesized through a structured LLM-driven pipeline consisting of content generation, segment-level annotation, and manual verification. Specifically, we adopt a two-stage prompting strategy to automatically construct intra-utterance emotion and duration TTS prompts. In Step 1, GPT-4o [^1] https://openai.com/index/hello-gpt-4o/ is prompted to generate 15,000 English and 15,000 Chinese emotion-rich content texts that explicitly exhibit continuous emotional transitions within a single utterance. Each text spans multiple emotional phases drawn from 7 core emotions (happy, sad, angry, surprised, fearful, disgusted, and neutral), forming either smooth or abrupt emotional progressions rather than a single static emotional state. In Step 2, DeepSeek-Chat  [^2] https://api-docs.deepseek.com/ is prompted to leverage its strong contextual understanding capability, performing precise semantic decomposition by segmenting the utterance into multiple emotion-specific segments. For each segment, it produces a concise natural language emotion description together with an estimated speaking duration, forming a structured sequence of emotion-duration pairs that directly aligns with the input requirements of controllable TTS systems. Example prompts used in Step 1 and Step 2 are shown in List. 8 and List. 9, respectively.
A.2 Quality Control and Manual Verification
At the automated level, we employ validation scripts to enforce strict structural, semantic, and distributional constraints. For Step 1 text generation, each sample is verified for valid JSON formatting and required fields, including the text content, text category, and emotion sequence. Specifically, text length is constrained to 15-25 words for English and 15-25 characters for Chinese texts, and each emotion sequence is restricted to 2-3 segments drawn from a predefined set of 7 emotion categories. To reduce redundancy, we remove exact duplicates and filter near-duplicate texts using similarity-based criteria computed from normalized sequence-matching scores between token sequences, applying an overall similarity threshold of 0.85 and an opening similarity threshold of 0.5 over the first few tokens, with similarity comparisons primarily performed among samples sharing the same emotion sequence. For Step 2 multi-segment prompt annotation, automated checks enforce strict alignment between the text and its associated emotion annotations. Specifically, the number and order of segments are required to exactly match the predefined emotion sequence. Each segment must include a valid emotion label, a non-empty natural language emotion description with constrained length (5-25 words for English or 5-30 characters for Chinese), and an estimated speaking duration falling within a predefined range of 0.3-8.0 seconds.
Beyond automated filtering, we conduct manual verification through stratified random sampling of the validated outputs. In total, 1,000 samples are randomly selected for human review, comprising 500 English and 500 Chinese samples. The sampling process is stratified by language, text category, and the number of emotion segments to ensure broad coverage across diverse data conditions. Human reviewers then examine segmentation boundaries, emotion-text alignment, vocal-affect descriptions, and the plausibility of estimated speaking durations to identify subtle issues that may escape rule-based automatic checks. Insights obtained from this process are used to iteratively refine prompting strategies and validation thresholds. The manual verification checklist applied in Step 1 and Step 2 is provided in List 10.
@@ A.3 Dataset Statistics and Distribution
(a) Distribution of Content Segments
(b) Distribution of Emotion Segments

### v2

# TED-TTS: Training-Free Intra-Utterance Emotion and Duration Control for Text-to-Speech Synthesis

@@ Abstract
While controllable Text-to-Speech (TTS) has achieved notable progress, most existing methods remain limited to inter-utterance-level control, making fine-grained intra-utterance expression challenging due to their reliance on non-public datasets or complex multi-stage training. In this paper, we propose TED-TTS, a training-free controllable framework for pretrained zero-shot TTS to enable intra-utterance emotion and duration expression. Specifically, we propose a segment-aware emotion conditioning strategy that combines causal masking with monotonic stream alignment filtering to isolate emotion conditioning and schedule mask transitions, enabling smooth intra-utterance emotion shifts while preserving global semantic coherence. Based on this, we further propose a segment-aware duration steering strategy to combine local duration embedding steering with global EOS logit modulation, allowing local duration adjustment while ensuring globally consistent termination. To eliminate the need for segment-level manual prompt engineering, we construct a 30,000-sample multi-emotion and duration-annotated text dataset to enable LLM-based automatic prompt construction. Extensive experiments demonstrate that our training-free method not only achieves state-of-the-art intra-utterance consistency in multi-emotion and duration control, but also maintains baseline-level speech quality of the underlying TTS model. Code [^1] Code: https://github.com/Simon-leong/TED-TTS. and audio samples [^2] Audio samples: https://simon-leong.github.io/TED-TTS-DemoPage/. are available.
TED-TTS: Training-Free Intra-Utterance Emotion and Duration Control for Text-to-Speech Synthesis
Qifan Liang1thanks: Equal contribution., Yuansen Liu [^1] footnotemark: 1, Ruixin Wei [^1] footnotemark: 1, Nan Lu1, Junchuan Zhao1, Ye Wang1 1School of Computing, National University of Singapore {liangqifan,yuansen,ruixin2003,nlu}@u.nus.edu, {junchuan,wangye}@comp.nus.edu.sg
@@ 1 Introduction
Humans naturally regulate emotional expression and speaking pace during speech in a dynamic and flexible manner, reflecting changes in semantics, emphasis, and discourse intent Scherer (2003); Tang et al. (2023). How to replicate such intra-utterance expressiveness remains a challenge in building human-like TTS synthesis systems.
@@
These challenges naturally raise an important question: Is it possible to achieve stable segment-level emotion transitions and duration control without retraining the model? In this paper, as shown in Fig. 1, we revisit controllable TTS from an inference-time perspective and propose the first Training-free Intra-Utterance Emotion and Duration control framework (TED-TTS). Rather than introducing additional predictors or retraining the acoustic model, our approach focuses on restructuring how conditioning information is accessed and updated during autoregressive decoding. Specifically, for multi-emotion control, we propose a segment-aware emotion conditioning strategy that combines causal masking with monotonic stream alignment algorithm, which jointly isolates segment-specific emotion conditioning and performs online text-semantic alignment to schedule mask transitions, enabling smooth intra-utterance emotion shifts while preserving global semantic coherence. To enable multi-duration control, we further propose a segment-aware duration steering strategy to incorporate local duration embedding steering with global EOS logit modulation, allowing segment-level pacing adjustment while ensuring globally consistent sequence termination. Besides, we construct a Multi-Emotion and Duration-annotated text dataset (MED-TTS) with 30,000 samples and fine-tune Qwen3-8B to enable LLM-based automatic prompt construction, thereby eliminating the need for segment-level manual segmentation and prompt engineering. Extensive experiments demonstrate that our method achieves state-of-the-art performance in stable intra-utterance multi-emotion transitions and duration control, while preserving the strong zero-shot synthesis capability of the underlying TTS model without any additional training. Our contributions are summarized as follows:
@@
  • We propose a segment-aware duration steering strategy that achieves local segment duration control while preserving globally consistent sequence termination.
@@ 3 Method
3.1 MED-TTS Dataset
Automatic Prompt Construction. Existing intra-utterance controllable TTS systems require manual text segmentation and segment-level emotion and duration specification, which is labor-intensive in real-world scenarios. To eliminate manual prompt engineering, we fine-tune the Qwen3-8B LLM to automatically transform raw user text into structured multi-segment prompts. As a prerequisite, we construct a dedicated Multi-Emotion and Duration-annotated text dataset (MED-TTS) with 30,000 samples, which is used to supervise emotion-aware text segmentation, natural language emotion description generation, and segment-level speech duration estimation. As illustrated below, MED-TTS is synthesized using LLM through a structured pipeline consisting of generation, annotation, and verification.
@@ 3.1 MED-TTS Dataset
Step 1: Content text generation. GPT-4o is prompted to generate 15,000 English and 15,000 Chinese emotion-rich content texts that explicitly exhibit continuous emotional transitions within a single utterance. Each text spans multiple emotional phases drawn from 7 core emotions (happy, sad, angry, surprised, fearful, disgusted, and neutral) and 3 text categories (descriptive, dialogue-style, and observational), forming either smooth or abrupt emotional progressions across diverse content genres. Example prompt used in Step 1 is shown in List. 9. Step 2: Multi-segment prompt annotation. To enable precise segment-level control, DeepSeek-Chat is prompted to perform precise semantic decomposition by segmenting the utterance into multiple emotion-specific segments. For each segment, it produces a concise natural language emotion description together with a realistic duration estimate, forming a structured sequence of emotion-duration pairs compatible with controllable TTS inputs. Example prompt used in Step 2 is shown in List. 10. Step 3: Post-processing and manual verification. Automatic checks are finally applied to filter samples with formatting errors, missing fields, or invalid segment boundaries, followed by systematic manual verification of outputs from both Step 1 and Step 2. The manual verification checklist is provided in List 11.
(a) Distribution of Content Segments
(b) Distribution of Emotion Segments
Dataset Statistics We summarize the distribution of the MED-TTS dataset across languages, text categories, and emotion segments. As illustrated in Fig. 3(a), the dataset is well balanced across languages, comprising 14,965 Chinese and 15,092 English samples. Within each language, samples are further evenly distributed across three text categories, each contributing approximately 5,000 utterances. Fig. 3(b) presents the segment-level emotion statistics. Across both Chinese and English, the 7 emotion types are uniformly represented, with each emotion accounting for approximately 12,000 segments. The average segment length remains stable within each language but differs across languages, with Chinese emotional segments typically spanning about 24-26 characters, while English segments are longer on average, ranging from roughly 34-39 words depending on emotion. Based on this dataset, we perform supervised fine-tuning with LoRA on the Qwen3-8B large language model, enabling automatic construction of segment-level TTS prompts without manual prompt engineering. Detailed prompting strategies, step-wise checklists, dataset statistics, and fine-tuning details are provided in the Appendix A.
@@ 3.2 Segment-Aware Emotion Conditioning
2D Causal Attention Mask. To resolve the misalignment between continuous generation and segment-level conditions, we design a 2D causal attention mask that disentangles condition visibility from semantic context. The mask preserves standard causal attention among text and semantic tokens across segment boundaries, ensuring globally coherent semantic generation while strictly restricting access to condition embeddings to be segment-local. Specifically, for any token that belongs to the $m$-th segment (either a text token in $\mathrm{X}_{m}$ or a generated semantic token that currently aligns to $\mathrm{X}_{m}$), attention is allowed to attend only to its corresponding condition embedding $\mathbf{C}_{m}$, while all other condition embeddings $\{\mathbf{C}_{j}\mid j\neq m\}$ are masked out. Meanwhile, each condition embedding $\mathbf{C}_{m}$ is prevented from attending to other condition embeddings, avoiding cross-condition information leakage. After that, as shown at the bottom of Fig. 2, emotional style is governed exclusively by the locally active condition, whereas semantic content remains globally visible through standard causal context.
@@ 5 Results and Evaluation
5.1 Comparison with Reference Models
5.2 Ablation Study
@@
Method $\times$0.75 $\times$0.875 $\times$1 $\times$1.125 $\times$1.25
@@
Monotonic Stream Alignment Evaluation. To evaluate the effectiveness of MSA, we visualize alignment results under different settings in Fig. 5 and report the mean absolute error (MAE) of segment boundary positions. Raw attention maps exhibit diffuse and locally non-monotonic patterns, making greedy alignment highly sensitive to noise and leading to unstable trajectories and frequent segment switching failures. Introducing the monotonic stream constraint alleviates this issue and reduces MAE to 0.216, but residual attention uncertainty still causes instability. By further incorporating the observation component, where the Top-$k$ most reliable attention heads are selected with $k=3$ instead of relying on a single maximum, MSA effectively suppresses alignment uncertainty, enforces smooth monotonic trajectories, and reduces MAE to 0.157, yielding precise emotion transitions closely aligned with the ground-truth boundaries.
Duration-Specified Evaluation. We evaluate duration-specified speech synthesis under five segment-level scaling factors ($\times$0.75, $\times$0.875, $\times$1.0, $\times$1.125, and $\times$1.25), comparing our full system with ablated variants and an IndexTTS2 baseline. As shown in Tab. 4, our method consistently achieves the lowest semantic token number error across all settings, reducing the error by 3.53% and 2.41% on average compared to variants without local steering and global EOS control, respectively. Relative to the baseline without explicit duration control, our approach further yields a 5.07% average error reduction, demonstrating accurate and robust duration control across diverse segment-level targets.
Figure 6: Efficiency analysis of the proposed framework under incremental configurations. S1 denotes the baseline model, S2 adds segment-aware emotion conditioning, and S3 further incorporates segment-aware duration steering. All metrics are evaluated under identical hardware conditions using the NVIDIA RTX 4090 GPU.
Inference Efficiency Evaluation. We evaluate the practical efficiency of the proposed framework using real-time factor (RTF), autoregressive (AR) latency, and end-to-end (E2E) latency, where latency is reported as the average inference time in seconds. Starting from the baseline (S1), we progressively incorporate segment-aware emotion conditioning (S2) with 2D causal masking and MSA, followed by segment-aware duration steering (S3) with local duration steering and global EOS steering. As shown in Fig. 6, incorporating emotion conditioning results in a moderate increase in computational cost, with RTF rising by 22.7% on English and 24.5% on Chinese. Further incorporating duration steering results in additional but controlled overhead, yielding total increases of 40.2% and 42.2%, respectively. Despite these increases, RTF remains consistently below 1.0 across both languages, indicating that real-time inference capability is preserved. Meanwhile, latency analysis shows that the majority of the overhead arises from the autoregressive stage. From S1 to S3, AR latency increases by 13.4% on English and 14.7% on Chinese, reflecting a favorable trade-off between controllability and efficiency while maintaining real-time performance.
@@ 6 Conclusion
In this paper, we propose the first training-free controllable framework to enable intra-utterance emotion and duration control in pretrained zero-shot TTS. By introducing segment-aware emotion conditioning and duration steering from an inference-time perspective, our method achieves smooth segment transitions within a single utterance, and is readily applicable to a broad class of autoregressive TTS backbones without requiring architectural modification or additional training. Extensive experiments demonstrate that our method not only delivers state-of-the-art intra-utterance controllability, but also preserves baseline-level speech quality of the underlying TTS model.
Acknowledgments
We thank the anonymous reviewers for their valuable comments and suggestions, as well as all participants involved in the subjective evaluations. Qifan Liang, Yuansen Liu, and Ruixin Wei contributed equally to this work. Junchuan Zhao is the corresponding author.
@@ A.1 LLM-Based Prompting Strategy
MED-TTS is synthesized through a structured LLM-driven pipeline consisting of content generation, segment-level annotation, and manual verification. Specifically, we adopt a two-stage prompting strategy to automatically construct intra-utterance emotion and duration TTS prompts. In Step 1, GPT-4o [^3] https://openai.com/index/hello-gpt-4o/ is prompted to generate 15,000 English and 15,000 Chinese emotion-rich content texts that explicitly exhibit continuous emotional transitions within a single utterance. Each text spans multiple emotional phases drawn from 7 core emotions (happy, sad, angry, surprised, fearful, disgusted, and neutral), forming either smooth or abrupt emotional progressions rather than a single static emotional state. In Step 2, DeepSeek-Chat  [^4] https://api-docs.deepseek.com/ is prompted to leverage its strong contextual understanding capability, performing precise semantic decomposition by segmenting the utterance into multiple emotion-specific segments. For each segment, it produces a concise natural language emotion description together with an estimated speaking duration, forming a structured sequence of emotion-duration pairs that directly aligns with the input requirements of controllable TTS systems. Example prompts used in Step 1 and Step 2 are shown in List. 9 and List. 10, respectively.
@@
A.2 Quality Control and Manual Verification
At the automated level, we employ validation scripts to enforce strict structural, semantic, and distributional constraints. For Step 1 text generation, each sample is verified for valid JSON formatting and required fields, including the text content, text category, and emotion sequence. Specifically, text length is constrained to 25-45 words for English and 20-30 characters for Chinese texts, and each emotion sequence is restricted to 2-3 segments drawn from a predefined set of 7 emotion categories. To reduce redundancy, we remove exact duplicates and filter near-duplicate texts using similarity-based criteria computed from normalized sequence-matching scores between token sequences, applying an overall similarity threshold of 0.85 and an opening similarity threshold of 0.5 over the first few tokens, with similarity comparisons primarily performed among samples sharing the same emotion sequence. For Step 2 multi-segment prompt annotation, automated checks enforce strict alignment between the text and its associated emotion annotations. Specifically, the number and order of segments are required to exactly match the predefined emotion sequence. Each segment must include a valid emotion label, a non-empty natural language emotion description with constrained length, and an estimated speaking duration falling within a predefined range of 0.3-8.0 seconds.
Beyond automated filtering, we conduct manual verification through stratified random sampling of the validated outputs. In total, 1,000 samples are randomly selected for human review, comprising 500 English and 500 Chinese samples. The sampling process is stratified by language, text category, and the number of emotion segments to ensure broad coverage across diverse data conditions. Human reviewers then examine segmentation boundaries, emotion-text alignment, vocal-affect descriptions, and the plausibility of estimated speaking durations to identify subtle issues that may escape rule-based automatic checks. Insights obtained from this process are used to iteratively refine prompting strategies and validation thresholds. The manual verification checklist applied in Step 1 and Step 2 is provided in List 11.